OpenStudy (anonymous):

The measure of a single angle in a regular octagon is (4x + 12)°. Find the value of x. Show equations and all work that leads to your answer.

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

First you have to find the measure of the angles in a rectangular octagon. See image. That should be easy. Once you've got that, just solve the equation.

OpenStudy (anonymous):

I got 1080 is the sum of the angles.

OpenStudy (anonymous):

Then i divided it by how many sides which there are 8 & i got 135 for the measure of each angle.

OpenStudy (anonymous):

the equation is 4x + 12 = 135?

OpenStudy (anonymous):

right?

OpenStudy (zehanz):

Right!

OpenStudy (zehanz):

How did you arrive at 1080?

OpenStudy (anonymous):

I used the formula (n-2)180

OpenStudy (anonymous):

so is x 8.44?

OpenStudy (anonymous):

@ZeHanz ??

OpenStudy (zehanz):

No, 4x=135-12=123, so x=123/4=30.75

OpenStudy (zehanz):

Sorry for the delay, I was also updating my iphone to iOS 6.1 and that went wrong :(

OpenStudy (anonymous):

oh alright.so each angle is = to 30.75?

OpenStudy (anonymous):

how do you check it to see if its right? @ZeHanz

OpenStudy (anonymous):

shouldn't 30.75 times 8 = the sum which is 1080?

OpenStudy (zehanz):

No, one angle is 4x+12, so to check do this: 8*(4*30.75+12) =...1080!